OpenStudy (anonymous):

Given the function y = 5x – 7, how does the y-value change as x increases by 1?

jimthompson5910 (jim_thompson5910):

pick any x value you want

jimthompson5910 (jim_thompson5910):

tell me which x value you picked

OpenStudy (anonymous):

The y-value decreases by 7?

jimthompson5910 (jim_thompson5910):

just pick any number you want and I'll show you how to find the answer

jimthompson5910 (jim_thompson5910):

Ok how about we pick the number 0, sound good?

OpenStudy (anonymous):

yah

jimthompson5910 (jim_thompson5910):

ok when x = 0 y = 5x – 7 y = 5(0) – 7 ... plug in x = 0 y = 0 - 7 y = -7 So when x = 0, y is y = -7

jimthompson5910 (jim_thompson5910):

So keep these two values in mind: x = 0 and y = -7

jimthompson5910 (jim_thompson5910):

Now bump x up by 1 to get x = 1 and repeat y = 5x – 7 y = 5(1) – 7 ... plug in x = 1 y = 5 - 7 y = -2 So when x = 1, y is y = -2

jimthompson5910 (jim_thompson5910):

Notice how when x goes from 0 to 1, y changes from -7 to -2 so as x goes up by 1, y goes up by 5 (the distance from -7 to -2 is 5 units)

jimthompson5910 (jim_thompson5910):

this is exactly what the slope is saying the slope is rise over run = 5/1 = (change in y of 5 units)/(change in x of 1 unit)

OpenStudy (anonymous):

ooooh so the y value increases by 5?

jimthompson5910 (jim_thompson5910):

exactly this is probably more evident if you use x values such as x = 3 and x = 4 When x = 3, y is... y = 5x – 7 y = 5(3) – 7 ... plug in x = 3 y = 15 - 7 y = 8 So when x = 3, y is y = 8 When x = 4, y is... y = 5x – 7 y = 5(4) – 7 ... plug in x = 4 y = 20 - 7 y = 13 So when x = 4, y is y = 13 The change from 8 to 13 is also 5 units, so that just further verifies the answer

OpenStudy (anonymous):

okayy thank you

jimthompson5910 (jim_thompson5910):

yw